Assume that a radiology practice has the following cost structure:

Fixed Costs = $400,000 Variable cost per procedure = $30

Furthermore, assume that the group expects to perform 7,000 procedures in the coming year.

  1. What is the group’s underlying cost structure?
  2. What is the group’s expected total variable costs?
  3. What is the group’s expected total costs?
  4. What is the group’s average cost per procedure?
Answer #1

What is the group’s underlying cost structure?

The underlying cos structure is a combination of fixed and variable cost.

What is the group’s expected total variable costs?

Total variable cost = Variable cost per procedure x Number of procedure = 30 x 7,000 = $ 210,000

What is the group’s expected total costs?

Total costs = Variable cost + fixed cost = 210,000 + 400,000 = $ 610,000

What is the group’s average cost per procedure?

Average cost = total cost / number of procedure = 610,000 / 7,000 = $ 87.14